Q.Choose the correct answer.
Q. The equation 2x2+7y2-4xy+3y-1=0 represents
a. A pair of straight lines
b. Ellipse
c. Parabola
d. Hyperbola
1
Expert's answer
2019-07-21T16:59:08-0400

I'll use invariants.



"2x^2+7y^2-4xy+3y-1=0 \\\\ \u00a0 \\\\\na_{11} = 2, a_{22} = 7, a_{12} = -2, a_{13} = 0, a_{23} = 1.5, a_{33} = -1 \\\\ \u00a0 \\\\I_2 = \\begin{vmatrix}\n 2 & -2 \\\\\n -2 & 7\n\\end{vmatrix} = 10 > 0 \\\\\u00a0\\\\I_3 = \\begin{vmatrix}\n 2 & -2 & 0 \\\\\n -2 & 7 & 1.5\\\\\n 0 & 1.5 & -1\n\\end{vmatrix} = -14 - 2*2.25 + 4 = -14.5 \\\\\u00a0\\\\ I_1 = tr\\begin{pmatrix}\n 2 & -2 \\\\\n -2 & 7\n\\end{pmatrix} = 9 \\\\\u00a0\\\\ I_2 > 0, \u00a0\u00a0 I_1I_3 < 0 \\implies Ellipse"


Answer: b